A particle is moving in the ~i - ~j plane with an acceleration of (1 - t)~ i+ t ~j where t is the time. The particle is projected from the point with an initial velocity of 3/2 ~i ms/1. Find the displacement of the particle at time t and the value of t when it is moving in the direction of motion.

What is the angle between its velocity and displacement at this time?


1
Expert's answer
2021-04-15T06:31:20-0400
"\vec a(t)=(1-t)\vec i+t\vec j"

"\vec v(t)=(t-\dfrac{1}{2}t^2)\vec i+\dfrac{1}{2}t^2 \vec j+\vec v_0"

"=(t-\dfrac{1}{2}t^2+\dfrac{3}{2})\vec i+\dfrac{1}{2}t^2 \vec j"

"\vec r(t)=(\dfrac{1}{2}t^2-\dfrac{1}{6}t^3+\dfrac{3}{2}t)\vec i+\dfrac{1}{6}t^3 \vec j+\vec r_0"

"\Delta \vec r=(\dfrac{1}{2}t^2-\dfrac{1}{6}t^3+\dfrac{3}{2}t)\vec i+\dfrac{1}{6}t^3 \vec j"


"\tan\theta=\dfrac{\dfrac{1}{6}t^3}{\dfrac{1}{2}t^2-\dfrac{1}{6}t^3+\dfrac{3}{2}t}"

"\tan\varphi=\dfrac{\dfrac{1}{2}t^2}{t-\dfrac{1}{2}t^2+\dfrac{3}{2}}"

Let "\tan\theta=\tan\varphi." Then


"\dfrac{\dfrac{1}{6}t^3}{\dfrac{1}{2}t^2-\dfrac{1}{6}t^3+\dfrac{3}{2}t}=\dfrac{\dfrac{1}{2}t^2}{t-\dfrac{1}{2}t^2+\dfrac{3}{2}}"

"2t^4-t^5+3t^3=3t^4-t^5+9t^3"


"t^4+6t^3=0"

Since "t>0," then there is no solution.


When particle is moving in the direction of motion the angle between its velocity and displacement is zero.